# Excluding cosmetic surgeries on hyperbolic 3-manifolds

### v2

@@ 1. Introduction
See Conjecture 1.3 for a version with multiple boundary tori, and Conjecture 1.6 for a version that seeks to exclude chirally cosmetic surgeries on knots in $S^{3}$. This paper provides extensive experimental evidence for all three conjectures.
@@ 1.1. Purely cosmetic surgeries on knots in $S^{3}$
Sometimes, these knot invariants can obstruct obstruct cosmetic surgeries on $N=S^{3}-K$ altogether. Notably, obstructions coming from knot Floer homology, combined with other invariants, have been used to prove Conjecture 1.2 for several infinite families of knots. These families include knots of genus $1$ (Wang [46]), torus knots (Ni and Wu [38]), two-bridge knots (Ichihara, Jong, Mattman, and Saito [27]), pretzel knots (Stipsicz and Szabo [39]), alternating knots with at least 7 twist regions (Ichihara and Jong [25]), connected sums (Tao [42]), and nontrivial cables (Tao [41]).
Knot invariants have also been used to verify Conjecture 1.2 on large data sets. The method is to first apply a test (such as one of the tests described in Section 2) to reduce the search to a small subset of knots, and to a finite and explicitly given set of slopes on these knots. Then, direct computation (often using hyperbolic invariants) can distinguish the Dehn fillings along the remaining slopes. In this vein, Hanselman used knot Floer homology (see Tests 2.72.9) to verify Conjecture 1.2 for all knots up to 16 crossings [18]. Detcherry used Jones polynomials (see Tests 2.52.6), with an assist from knot Floer homology, to verify Conjecture 1.2 for all knots up to 17 crossings [10].
@@
The method of proof is still to compute obstructions coming from knot invariants, reducing the sample dramatically. One innovation, compared to prior work, is that direct computation of knot Floer homology is never needed for the knots in the set. Instead, we rely heavily on the Alexander polynomial and the Turaev genus (see Test 2.4), to capture most of the obstructive information contained in knot Floer homology. This speeds up computations dramatically, making it feasible to check the 352,152,252 prime knots up to 19 crossings.
@@ 1.2. Cosmetic surgery conjecture for general 3-manifolds
[…]
@@
Conjectures 1.2 and 1.3 hold for the 59,107 one-cusped manifolds in the SnapPy census, provided that SnapPy’s systole calculation is correct to within 1%.
The proof of Theorem 4.4 makes heavy use of hyperbolic geometry. For each manifold $N$, we wish to restrict attention to a small, explicit list of slopes. Thus we begin by proving Conjecture 1.3 for all sufficiently long (tuples of) slopes, where “long” is explicitly quantified. For these long slopes, the only purely or chirally cosmetic surgeries come from symmetries of $N$ itself. An ineffective version of this sort of statement appears in Bleiler–Hodgson–Weeks [3], and an effective version (with an explicit quantification) appears in our previous paper [15, Theorem 7.29].
@@
When $N$ has more than one cusp, let ${\mathbf{s}}=(s_{1},\ldots,s_{n})$ be a tuple of slopes, with one slope per cusp. The normalized length of ${\mathbf{s}}$ is the quantity $L({\mathbf{s}})$ satisfying $1/L({\mathbf{s}})^{2}=\sum_{i=1}^{n}1/L(s_{i})^{2}$.
@@
Theorem 3.19 is a slight strengthening of [15, Theorem 7.29]. Compared to that statement, both the constant term and the variable term in the maximum of (1.5) have become slightly smaller. This makes Theorem 3.19 easier to apply in practice.
In Section 4, we explain how Theorem 3.19 leads to a practical procedure to check the Cosmetic surgery conjecture on one-cusped manifolds. Here is an outline. When $N$ has one cusp, there are only finitely many slopes $s$ whose normalized length is shorter than the bound in (1.5). Supposing for concreteness that ${\operatorname{sys}}(N)\geq 0.145$, the length bound becomes $9.97$. By a lemma of Agol [1, Lemma 8.2], there are at most $102$ slopes on the cusp of $N$ of normalized length at most $10$, hence $s$ must be one of these 102 slopes. If $N(s)\cong N(s^{\prime})$ and this manifold is hyperbolic, then volume considerations imply that $L(s^{\prime})$ is also bounded above; see Theorem 3.24. This leads to an explicit finite list of slope pairs $(s,s^{\prime})$ to check, and every pair of filled manifolds $N(s),N(s^{\prime})$ can be compared by computer. The finite list of pairs of exceptional surgeries on $N$ can also be compared by computer.
We have implemented this procedure in Sage, and the program is publicly available on GitHub [14]. For any given one-cusped manifold $N$, the program forms the finite set of pairs of slopes that need to be checked, and then proceeds to compare them. Section 4 describes the procedure in detail. All of the comparisons between manifolds are done using rigorously verified hyperbolic invariants, or else using algebraic invariants such as the homology of covers. However, the computation of ${\operatorname{sys}}(N)$ is a crucial input for building the set of slopes to check, and this computation is not currently rigorously verified in SnapPy. This is why Theorem 4.4 contains the proviso that SnapPy’s systole calculation is correct to within $1\%$. See Remark 4.3 for more discussion.
@@ 1.3. Chirally cosmetic surgeries
There are no known counterexamples to Conjecture 1.2, that is, no known examples of purely cosmetic surgeries on a 3-manifold $N$ with incompressible boundary However, as stated above, there exist known examples of chirally cosmetic surgeries. Recently Ichihara, Ito, and Saito proposed the following conjecture [23, Conjecture 1], which states that chirally cosmetic surgeries on knots in the 3-sphere only occur in very restricted ways.
@@
See [23, Conjecture 1] for a slightly stronger formulation: chirally cosmetic surgeries on knots in $S^{3}$ either satisfy Conjecture 1.6 or are particular pairs of slopes on $(2,r)$ torus knots. We restrict attention to non-torus knots because our methods are primarily hyperbolic.
[…]
@@
Conjecture 1.6, the chirally cosmetic surgery conjecture for knots in $S^{3}$, holds for the 313,209 hyperbolic knots up to 15 crossings, provided that SnapPy’s systole calculation is correct to within 1%.
In prior work, Ichihara, Ito, and Saito have verified Conjecture 1.6 for approximately 75% of the 245 non-$(2,r)$-torus knots up to 10 crossings [23, Table 2]. Theorem 4.7 expands this verification to a larger data set, contingent on the accuracy of the systole calculation. The proviso about systoles occurs for the same reason as in Theorem 4.4.
@@ 1.4. Common fillings of distinct cusped manifolds
In Section 5, we walk through the adaptations needed to consider distinct parent 3-manifolds. We apply this in two results. In Theorem 5.1, we find all knots with at most 13 crossings that share a common Dehn surgery with the figure-8 knot. In Theorem 5.2, we find all knot complements that can be triangulated with fewer than 10 tetrahedra that share a common Dehn filling with the figure-8 knot complement. These results have been used by Kalfagianni and Melby [32] to construct many new $q$-hyperbolic knots in $S^{3}$.
@@ 1.5. Code on GitHub
At the same time, we have not implemented a complete solution to the homeomorphism problem for 3-manifolds. Accordingly, our programs are not guaranteed to distinguish every non-homeomorphic pair of Dehn fillings of $N$. We have entered a number of invariants — hyperbolic volume, the homology of covers, Seifert invariants — but we have necessarily had to stop somewhere. The proof of Theorem 4.4 contains several examples of toroidal fillings that had to be handled “by hand.” We emphasize that, apart from systole calculations, all of the tests used by our programs are rigorous.
We point the reader to Remarks 2.13, 4.5, 4.8, and 5.3 for a description of the top-level routines that check the conjectures, as well as for some notes about running time.
@@ 1.6. Acknowledgements
As the computational and theoretical aspects of this project spanned several years, we have had occasion to lean on the input many mathematicians. We thank Steve Boyer for suggesting that the Casson invariant works in settings more general than $S^{3}$. We thank Jonathan Hanselman for explaining his work and its many applications. We thank Renaud Detcherry, Kazuhiro Ichihara, Tetsuya Ito, and Toshio Saito for patiently answering questions about quantum invariants, especially in the setting of chirally cosmetic surgeries.
@@
We thank Adam Lowrance for sharing his code for computing the diagram Turaev genus. We thank Nathan Dunfield and Giles Gardam for suggesting the use of normal cores of subgroups (Definition 3.2) as a computable invariant of fundamental groups. We also thank Dunfield for writing code tying together SnapPy and Regina [12], for answering numerous computational questions, and for upgrading several SnapPy features (especially cover enumeration) in a way that proved very timely for this project.
@@
Futer was partially supported by NSF grant DMS–1907708. Purcell was partially supported by ARC grant DP210103136. We thank Temple University and the Oberwolfach Mathematics Institute for their hospitality during collaborative visits. All of the computations described in this paper were carried out on a 2021 M1 iMac, purchased for this project using NSF grant DMS–1907708.
@@ 2. Obstructions and tests from knot invariants
Several of the tests below rely on a recent theorem of Hanselman [18, Theorem 2]. In the following theorem, $g(K)$ is the Seifert genus of a knot $K$. Meanwhile, $th(K)$ is an invariant derived from the knot Floer homology $\widehat{HFK}(K)$. Recall that $\widehat{HFK}(K)$ is a bigraded, finitely generated abelian group, whose nontrivial entries lie on finitely many slope–$1$ lines with respect to the bigrading. The thickness $th(K)$ is the largest difference between the $y$–intercepts of these lines.
@@ 2. $\>\>(2)$
In particular, if $th(K)<2g(K)(g(K)-2)$, then $K$ cannot have purely cosmetic surgeries.
Theorem 2.1 is a powerful result that inspired a considerable amount of subsequent work. Using it directly, as we do in Test 2.9, is a costly computation because the number of generators in the chain complex for $\widehat{HFK}(K)$ is factorial in the crossing number $c(K)$. Fortunately, both the thickness $th(K)$ and the genus $g(K)$ can be estimated using simpler invariants.
@@
Given a knot diagram $D=D(K)$, the diagram Turaev genus $g_{T}(D)$ is an invariant of $D$ that can be computed from a single Kauffman state of $D$. See [6, Section 1.1] for the exact definition. The Turaev genus of $K$, denoted $g_{T}(K)$, is a knot invariant defined by minimizing $g_{T}(D)$ over all diagrams of $K$. A knot $K$ is alternating if and only if $g_{T}(K)=0$ [9, Corollary 4.6].
@@
The following tests are ordered roughly according to computational difficulty. The first two tests use the Alexander polynomial $\Delta_{K}(t)$.
Test 2.3 (Casson invariant test).
Given a 3-manifold $N=Y-K$, where $Y$ is an integer homology sphere, compute the (symmetrized) Alexander polynomial $\Delta_{N}=\Delta_{K}$ and the Casson invariant $\frac{1}{2}\Delta^{\prime\prime}_{K}(1)$. If $\Delta^{\prime\prime}_{K}(1)\neq 0$, then a theorem of Boyer and Lines [4, Proposition 5.1] implies that $N$ cannot have purely cosmetic surgeries.
We remark that Test 2.3 is the only test described in this section that works outside the realm of knot complements in $S^{3}$.
Test 2.4 (Quick genus–thickness test).

### v3

@@ 1. Introduction
This problem appears as [21, Conjecture 6.1], as well as [3, Problem 1.9(a)]. See also Conjecture 1.3 for a version with multiple boundary tori, and Conjecture 1.6 for a version that seeks to exclude chirally cosmetic surgeries on knots in $S^{3}$. This paper provides extensive experimental evidence for all three conjectures.
@@ 1.1. Purely cosmetic surgeries on knots in $S^{3}$
Sometimes, these knot invariants can obstruct cosmetic surgeries on $N=S^{3}-K$ altogether. Notably, obstructions coming from Heegaard Floer homology, combined with other invariants, have been used to prove Conjecture 1.2 for several infinite families of knots. These families include knots of genus $1$ (Wang [51]), torus knots (Ni and Wu [43]), two-bridge knots (Ichihara, Jong, Mattman, and Saito [32]), pretzel knots (Stipsicz and Szabo [44]), alternating knots with at least 7 twist regions (Ichihara and Jong [30]), connected sums (Tao [47]), and nontrivial cables (Tao [46]).
After this paper was first distributed, Daemi, Lidman, and Miller Eismeier used instanton Floer homology to prove that the Alexander polynomial obstructs cosmetic surgeries [10]; see Theorem 2.2 for details. It follows as a consequence that Conjecture 1.2 holds for all alternating knots, all fibered knots, and all knots of genus $g\neq 2$.
Knot invariants have also been used to verify Conjecture 1.2 on large data sets. The method is to first apply a test (such as one of the tests described in Section 2) to reduce the search to a small subset of knots, and to a finite and explicitly given set of slopes on these knots. Then, direct computation (often using hyperbolic invariants) can distinguish the Dehn fillings along the remaining slopes. In this vein, Hanselman used knot Floer homology (see Tests 2.92.10) to verify Conjecture 1.2 for all knots up to 16 crossings [23]. Detcherry used Jones polynomials (see Tests 2.72.8), with an assist from knot Floer homology, to verify Conjecture 1.2 for all knots up to 17 crossings [12].
@@
The method of proof is still to compute obstructions coming from knot invariants, reducing the sample dramatically. One innovation, compared to prior work, is that direct computation of knot Floer homology is never needed for the knots in the set. Instead, we rely heavily on the Alexander polynomial and the Turaev genus (see Test 2.6), to capture most of the obstructive information contained in knot Floer homology. This speeds up computations dramatically, making it feasible to check the 352,152,252 prime knots up to 19 crossings.
@@ 1.2. Cosmetic surgery conjecture for general 3-manifolds
Compare Kirby [38, Problem 1.81(B)] for a closely related statement, which is too strong to be true. Indeed, restricting to purely cosmetic surgeries is necessary, even when $N$ has a single cusp. For instance, Dunfield (personal communication) has observed that in the census manifold $N=\texttt{o9\_39009}$, the slopes (1,3)(-1,3) and (3,2)(-3,2) are chirally cosmetic, with the further property that the core curves of Dehn filling solid tori are isotopic to geodesics. Ichihara and Jong [31] have also constructed a one-cusped hyperbolic manifold admitting chirally cosmetic surgeries.
@@
Conjectures 1.2 and 1.3 hold for the 59,107 one-cusped manifolds in the SnapPy census.
The proof of Theorem 4.3 makes heavy use of hyperbolic geometry. For each manifold $N$, we wish to restrict attention to a small, explicit list of slopes. Thus we begin by proving Conjecture 1.3 for all sufficiently long (tuples of) slopes, where “long” is explicitly quantified. For these long slopes, the only purely or chirally cosmetic surgeries come from symmetries of $N$ itself. An ineffective version of this sort of statement appears in Bleiler–Hodgson–Weeks [4], and an effective version (with an explicit quantification) appears in our previous paper [18, Theorem 7.29].
@@
When $N$ has multiple cusps, let ${\mathbf{s}}=(s_{1},\ldots,s_{n})$ be a tuple of slopes, with one slope per cusp. The normalized length of $\mathbf{s}$} is the quantity $L({\mathbf{s}})$ satisfying $1/L({\mathbf{s}})^{2}=\sum_{i=1}^{n}1/L(s_{i})^{2}$.
@@
Theorem 3.19 is a slight strengthening of [18, Theorem 7.29]. Compared to that statement, both the constant term and the variable term in the maximum of (1.5) have become slightly smaller. This makes Theorem 3.19 easier to apply in practice.
In Section 4, we explain how Theorem 3.19 leads to a practical procedure to check the Cosmetic surgery conjecture on one-cusped manifolds. Here is an outline. When $N$ has one cusp, there are only finitely many slopes $s$ whose normalized length is shorter than the bound in (1.5). Suppose for concreteness that ${\operatorname{sys}}(N)\geq 0.145$, hence the length bound becomes $9.97$. By a lemma of Agol [1, Lemma 8.2], there are at most $102$ slopes on the cusp of $N$ of normalized length at most $10$, hence $s$ must be one of these 102 slopes. If $N(s)\cong N(s^{\prime})$ and this manifold is hyperbolic, then volume considerations imply that $L(s^{\prime})$ is also bounded above; see Theorem 3.24. This leads to an explicit finite list of slope pairs $(s,s^{\prime})$ to check, and every pair of filled manifolds $N(s),N(s^{\prime})$ can be compared by computer. The finite list of pairs of exceptional surgeries on $N$ can also be compared by computer.
We have implemented this procedure in Sage, and the program is publicly available on GitHub [17]. For any given one-cusped manifold $N$, the program forms the finite set of pairs of slopes that need to be checked, and then proceeds to compare them. Section 4 describes the procedure in detail. All of the comparisons between manifolds are done using rigorously verified hyperbolic invariants, or else using algebraic invariants such as the homology of covers. This includes the rigorous verification of ${\operatorname{sys}}(N)$, which is included in SnapPy versions 3.2 and following. The forthcoming work of Goerner, Haraway, Hoffman, and Trnkova [20] includes an explanation of the verified length spectrum algorithm and a justification of the error bounds that are used in SnapPy 3.3.
As with all computer-assisted proofs, the correctness relies upon the rigor of our mathematical arguments (described in Section 3), of our code (described in Section 4), and of the software packages that we use (Sage, SnapPy, and Regina).
@@ 1.3. Chirally cosmetic surgeries
There are no known counterexamples to Conjecture 1.2, that is, no known examples of purely cosmetic surgeries on a 3-manifold $N$ with incompressible boundary However, as stated above, there exist known examples of chirally cosmetic surgeries. Recently Ichihara, Ito, and Saito proposed the following conjecture [28, Conjecture 1], which states that chirally cosmetic surgeries on knots in the 3-sphere only occur in very restricted ways.
@@
See [28, Conjecture 1] and [3, Propblem 1.9(d)] for a slightly stronger formulation: chirally cosmetic surgeries on knots in $S^{3}$ either satisfy Conjecture 1.6 or are particular pairs of slopes on $(2,r)$ torus knots. We restrict attention to non-torus knots because our methods are primarily hyperbolic.
The statements of Conjecture 1.6 and of [28, Conjecture 1] are carefully designed to get around known counterexamples. For instance, $(2,r)$ torus knots are known to admit chirally cosmetic surgeries; see [29, Corollary A.2]. There are also known examples of one-cusped hyperbolic manifolds that do not embed in $S^{3}$ and admit chirally cosmetic surgeries, such as $N=\texttt{o9\_39009}$ and the example of Ichihara and Jong [31].
@@
Conjecture 1.6, the chirally cosmetic surgery conjecture for knots in $S^{3}$, holds for the 313,209 hyperbolic knots up to 15 crossings.
In prior work, Ichihara, Ito, and Saito have verified Conjecture 1.6 for approximately 75% of the 245 non-$(2,r)$-torus knots up to 10 crossings [28, Table 2]. Theorem 4.6 expands this verification to a much larger data set.
@@ 1.4. Common fillings of distinct cusped manifolds
In Section 5, we walk through the adaptations needed to consider distinct parent 3-manifolds. We apply this in two results. In Theorem 5.1, we find all knots with at most 13 crossings that share a common nontrivial Dehn surgery with the figure-8 knot. In Theorem 5.2, we find all knot complements that can be triangulated with fewer than 10 tetrahedra that share a common nontrivial Dehn filling with the figure-8 knot complement. These results have been used by Kalfagianni and Melby [37] to construct many new $q$-hyperbolic knots in $S^{3}$.
@@ 1.5. Code on GitHub
At the same time, we have not implemented a complete solution to the homeomorphism problem for 3-manifolds. Accordingly, our programs are not guaranteed to distinguish every non-homeomorphic pair of Dehn fillings of $N$. We have entered a number of invariants — hyperbolic volume, length spectra, the homology of covers, Seifert invariants — but we have necessarily had to stop somewhere. Indeed, the proof of Theorem 4.3 contains several examples of Dehn fillings that had to be handled “by hand.” We emphasize that all of the tests used by our programs are rigorous.
We point the reader to Remarks 2.14, 4.4, 4.7, and 5.3 for a description of the top-level routines that check the conjectures, as well as for some notes about running time.
@@ 1.6. Acknowledgements
As the computational and theoretical aspects of this project spanned several years, we have had occasion to lean on the input of many mathematicians. We thank Steve Boyer for suggesting that the Casson invariant works in settings more general than $S^{3}$. We thank Jonathan Hanselman and Tye Lidman for explaining their work, including its many applications. We thank Renaud Detcherry, Kazuhiro Ichihara, Tetsuya Ito, and Toshio Saito for patiently answering questions about quantum invariants, especially in the setting of chirally cosmetic surgeries.
@@
We thank Nathan Dunfield and Matthias Goerner for answering numerous computational questions, and for implementing several SnapPy features (fast cover enumeration and verified length spectra) in a way that proved very timely for this project. We also thank Dunfield for writing code tying together SnapPy and Regina [13]. We thank Adam Lowrance for sharing his code for computing the diagram Turaev genus. We thank both Nathan Dunfield and Giles Gardam for suggesting the use of normal cores of subgroups (Definition 3.2) as a computable invariant of fundamental groups.
@@
Futer was partially supported by NSF grants DMS–1907708 and DMS–2405046. Purcell was partially supported by ARC grant DP210103136. We thank Temple University and the Oberwolfach Mathematics Institute for their hospitality during collaborative visits. All of the computations described in this paper were carried out on a 2021 M1 iMac, purchased for this project using NSF grant DMS–1907708.
@@ 2. Obstructions and tests from knot invariants
2.1. Theoretical tools
Several of the tests below rely on a recent theorem of Hanselman [23, Theorem 2]. In the following theorem, $g(K)$ is the Seifert genus of a knot $K$. Meanwhile, $th(K)$ is an invariant derived from the knot Floer homology $\widehat{HFK}(K)$. Recall that $\widehat{HFK}(K)$ is a bigraded, finitely generated abelian group, whose nontrivial entries lie on finitely many lines of slope $1$ with respect to the bigrading. The thickness $th(K)$ is the largest difference between the yy–intercepts of these lines.
@@ 2.1. Theoretical tools
In particular, if th(K)<2g(K)(g(K)2)th(K)<2g(K)(g(K)-2), then $K$ cannot have purely cosmetic surgeries.
Much more recently, after this paper was first distributed, Daemi, Lidman, and Miller Eismeier proved a powerful theorem [10, Theorems 1.2 and 6.1] that restricts the possibilities for cosmetic surgeries even further.
Theorem 2.2 (Daemi, Lidman, Miller Eismeier).
Let $K\subset S^{3}$ be a nontrivial knot. Then
  1. (1)\>\>(1)
    The pair of slopes $\{\pm\tfrac{1}{q}\}$ cannot be purely cosmetic for $K$.
  2. (2)\>\>(2)
    If $\{\pm 2\}$ is a purely cosmetic pair for $K$, then the Alexander polynomial of $K$ is $\Delta_{K}=1$.
Combining Theorems 2.1 and 2.2, we learn that any nontrivial knot in $S^{3}$ admitting a purely cosmetic pair of slopes $s_{1},s_{2}$ must have $\Delta_{K}=1$. This leads to Test 2.4, which is both powerful and easy to apply. It also implies that Conjecture 1.2 holds for large families, including alternating knots and fibered knots.
Even taken alone, Theorem 2.1 is a powerful result that inspired a considerable amount of subsequent work. Using it directly, as we do in Test 2.10, is a costly computation because the number of generators in the chain complex for $\widehat{HFK}(K)$ is factorial in the crossing number $c(K)$. Fortunately, both the thickness $th(K)$ and the genus $g(K)$ can be estimated using simpler invariants.
@@
Given a knot diagram $D=D(K)$, the diagram Turaev genus $g_{T}(D)$ is an invariant of $D$ that can be computed from a single Kauffman state of $D$. See [7, Section 1.1] for the exact definition. The Turaev genus of $K$, denoted $g_{T}(K)$, is a knot invariant defined by minimizing $g_{T}(D)$ over all diagrams of $K$. A knot $K$ is alternating if and only if $g_{T}(K)=0$ [11, Corollary 4.6].
@@
2.2. Practical tests
The following tests are ordered roughly according to computational difficulty. The first three tests use the Alexander polynomial $\Delta_{K}(t)$, normalized to be symmetric in $t$ and $t^{-1}$.
Test 2.4 (Alexander polynomial test).
Given a knot $K\subset S^{3}$, compute the Alexander polynomial $\Delta_{K}$. If $\Delta_{K}\neq 1$, then Theorems 2.1 and 2.2 imply that $K$ cannot have purely cosmetic surgeries.
Test 2.5 (Casson invariant test).
Given a 3-manifold N=YKN=Y-K, where $Y$ is an integer homology sphere, compute the (symmetrized) Alexander polynomial $\Delta_{N}=\Delta_{K}$ and the Casson invariant $\frac{1}{2}\Delta^{\prime\prime}_{K}(1)$. If $\Delta^{\prime\prime}_{K}(1)\neq 0$, then a theorem of Boyer and Lines [5, Proposition 5.1] implies that $N$ cannot have purely cosmetic surgeries.
We remark that Test 2.5 is the only test described in this section that works outside the realm of knot complements in $S^{3}$. For this reason, it is not strictly subsumed by Test 2.4.
Test 2.6 (Quick genus–thickness test).